3.1 There are two countries, the US and Canada. The two countries produce tomatoes and corn. A US worker can produce 200 pounds of tomatoes or 50 bushels of corn per year. The US has 300 workers. Canadian workers can produce 160 pounds of tomatoes or 20 bushel of corn per year. Canada has 30 workers.

a. Which country has an absolute advantage in produce tomatoes? In corn? (1 point)

b. Which country has a comparative advantage in produce tomatoes? In corn? (2 points)

c. Draw the US’s production possibilities curve (placing corn on the x-axis). What is the relative price of corn in the US? (2 points)

d. Draw Canada’s production possibilities curve (placing corn on the x-axis). What is the relative price of corn in Canada? (2 points)

e. Assuming a relative price PC/PT = 6 lbs of tomatoes per bushel of corn, draw the trade line for the US and show how the US can gain from trade. Use the graph in part c. (3 points)
Answer #1

(a)

Each US worker can produce more tomato than in Canada (200 > 160) and more corn than in Canada (50 > 20). So US has absolute advantage in both tomato and corn.

(b)

We compute opportunity cost (OC) as follows.

In US,

OC of tomato = 50/200 = 0.25 corn

OC of corn = 200/50 = 4 tomato

In Canada,

OC of tomato = 20/160 = 0.125 corn

OC of corn = 160/20 = 8 tomato

Since Canada can produce tomato at lower OC than US can (0.125 < 0.25), Canada has comparative advantage in tomato.

Accordingly, US has comparative advantage in corn.

(c) and (d):

Output table for the countries as follows.

US (300 workers) Canada (30 workers)
Tomato 300 x 200 = 60,000 30 x 160 = 4,800
Corn 300 x 50 = 15,000 30 x 20 = 600

(c)

PPF of US as follows.

Relative price of corn in US = OC of corn = 8 tomato

(d)

PPF of Canada as follows.

Relative price of corn in Canada = OC of corn = 4 tomato

NOTE: As per Chegg Answering Policy, 1st 4 parts have been answered.